OpenStudy (anonymous):

Given the function f(x) = 5(x+4) − 6, solve for the inverse function when x = 19. SOMEONE HELP ME I KEEP GETTING 109 and that's not an answer .

OpenStudy (anonymous):

solve \[5(x+4)-6=19\]

OpenStudy (anonymous):

add 6 to get \[5(x+4)=25\] divide by 5 and get \[x+4=5\]

OpenStudy (anonymous):

but if x = 19 wouldn't you substitute both and get 19= 5(19+4)-6

OpenStudy (anonymous):

I got 109